OpenStudy (anonymous):

Tom receives allowance every week. He spends three-fourths of his allowance on food. Last week, Tom spent nine dollars on food. What is Tom’s weekly allowance? Write the verbal model, define the variable, write the equation and find the solution. Show your work.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

3/4 of x = 9 3/4x = 9 x = 9/(3/4) x = 9 * 4/3 x = 36/3 = 12

OpenStudy (anonymous):

understand ?

OpenStudy (anonymous):

3/4 of x = $9 or, 3/4 times x = $9 or, 3x/4 = $9 or, 3x = $9*4 ( by shifting 4 to right side) or, 3x = $36 ( 9 times 4 equals 36) or, x = $36/3 ( by shifting 3 to right side) or, x = $ 12 ( 36 divides b6 3 equals 12) so, his weekly allowance is $12. yes is this correct

OpenStudy (anonymous):

yes..his weekly allowance is $12

OpenStudy (anonymous):

ok